OpenStudy (anonymous):

a triangle has an area of 85 cm2. if two of the sides are 18 and 11 what is the third side?

OpenStudy (astrophysics):

\[A = 1/2 bh \] 

OpenStudy (astrophysics):



OpenStudy (anonymous):

...

OpenStudy (anonymous):

a can't be equal to 11

OpenStudy (javk):

Let me know what you get

OpenStudy (wolf1728):

If area = 85 and area = .5 * b * h If base = 18 then height = 85 / (18*.5) height = 9.4444444444 Using Heron's formula, area = square root (s • (s - a) • (s - b) • (s - c)) where s = semi perimeter (lengths of all 3 sides divided by 2) area^2 = s • (s - a) • (s - b) • (s - c) 85^2 = s • (s - 18) • (s - 11) • (s - c) 7,225 = s^2 - 18s * s^2 -11s -cs +11c If you could solve for 'c' it would equal 15.5555555555555

OpenStudy (astrophysics):

a is 11, b represents the whole line segment.

OpenStudy (adi3):

wait guys pause for a sec srry, is tht higher level maths

OpenStudy (wolf1728):

Use Heron's formula sides = 18, 11 and 15.555555555555555555555 s = semi-perimeter = (18 + 11 + 15.5555555555)/2 s = 22.2777777778 area = square root (s • (s - a) • (s - b) • (s - c)) area = square root (22.2777777778 * (22.2777777778 -18) * (22.2777777778-11) * (22.2777777778 -15.55555555555) area = sq root (22.2777777778 * * 4.2777777778 * 11.2777777778 * 6.7222222222) area = sq root (7,224.8109186922) area = 84.9988877497

OpenStudy (javk):

Actually its really simple stuff @Adi3, its just simple manipulation of the Pythagoras theorem and the area of a triangle with a little bit of imagination and visualization :) I find it helpful to form a quick mental picture of what's going on before I start, and if need be I might jot it down on a piece of paper. What do you say @wolf1728?

OpenStudy (anonymous):

15.56 is my answer and yes it's correct thankss.. i was actually esay i just drew it incorrectly . :D :D thank you all especially @Javk